# LaM-SLidE: Latent Space Modeling of Spatial Dynamical Systems via Linked Entities

### v3

@@ 1 Introduction
A conventional approach to predict spatial trajectories of entities is to represent scenes as neighborhood graphs and to subsequently process these graphs with graph neural networks (GNNs). When using GNNs [83, 64, 32, 11], each entity is usually represented by a node, and the spatial entities nearby are connected by an edge in the neighborhood graph. Neighborhood graphs have extensively been used for trajectory prediction tasks [51], especially for problems with a large number of indistinguishable entities, [e.g., 81, 63]. Recently, GNNs have been integrated into generative modeling frameworks to effectively capture the behavior of stochastic systems [97, 22].
Despite their widespread use in modeling spatial trajectories, GNNs hardly follow recent trends in latent space modeling, where unified representations [45] together with universality and scalability of transformer blocks [90] offer simple application across datasets and tasks, a behavior commonly observed in computer vision and language processing [24, 25]. Notably, recent breakthroughs in image and video generation can be accounted to latent space generative modeling [38, 15]. In such paradigms, pre-trained encoders and decoders are employed to map data into a latent space, where subsequent modeling is performed, leveraging the efficiency and expressiveness of this representation. This poses the question:
Can we leverage recent techniques from generative latent space modeling to boost the modeling of stochastic trajectories of dynamical systems with varying number of entities?
@@
Figure 1: Overview of our approach. Left: Conventional graph neural networks (GNNs) model time-evolving systems (e.g., molecular dynamics) by representing entities as nodes and iteratively updating node embeddings and positions to capture system dynamics across timesteps. Right: Latent diffusion models employ an encoder-decoder architecture to compress input data into a lower-dimensional latent space where generative modeling is performed. Latent diffusion models, frequently enhanced with conditional information like text, excel at generative tasks, but due to the fixed input/output structure, are not directly adaptable to physical systems with varying number of entities. Middle: Our proposed approach LaM-SLidE bridges these paradigms by: (1) introducing identifiers that allow traceability of individual entities, and (2) leveraging a latent system representation.
In order to leverage a latent systems representation, we need to be able to trace individual entities of the system. The core idea of LaM-SLidE is the introduction of identifiers (IDs) that allow for retrieval of entity properties, e.g. entity coordinates, from the latent system representation. Consequently, we can train generative models, like flow-matching [54, 57, 2], purely in latent space, where pre-trained decoder blocks map the generated representations back to the physics domain. An overview is given in Figure 1. Qualitatively, LaM-SLidE demonstrates flexibility and perform favorable across a variety of different modeling tasks.
@@
  • Cross-domain generalization: We perform experiments in different domains with varying degrees of difficulty, focusing on molecular dynamics, human motion behavior, and particle systems. LaM-SLidE performs favorably with respect to all other architectures and showcasing scalability with model size.
@@ 2 Background & Related Work
We note that $\Phi$ does not necessarily need to be defined on the whole space ×$\mathbb{R}\times\mathcal{S}$, but we assume this for notational simplicity. The exact formal definition of random dynamical systems is more involved and consists of a base flow (noise) and a cocycle dynamical system defined on a physical phase space [8]. We skip the details, but assume to deal with random dynamical systems for the remainder of the paper. The non-deterministic behavior of such dynamical systems suggests generative modeling approaches.
@@
Recent developments in generative modeling have captured widespread interest. The breakthroughs of the last years were mainly driven by diffusion models [86, 87, 37], a paradigm that transforms a simple distribution into a target data distribution via iterative refinement steps. Flow Matching [54, 57, 2] has emerged as a powerful alternative to diffusion models, enable simulation-free training between arbitrary start and target distributions [55] and were also extended to data manifolds [19]. This approach comes with straighter paths, offering faster integration, and has been successfully applied across different domains like images [27], audio [91], videos [71], protein design [41] and robotics [13].
@@
Latent space modeling has achieved remarkable success at image and video generation [15, 27], where pre-trained encoders and decoders map data into a latent space, and back into the physics space. The latent space aims to preserve the essential structure and features of the original data, often following a compositional structure $\mathcal{D}\circ\mathcal{A}\circ\mathcal{E}$ [85, 4, 5], where the encoder $\mathcal{E}$ maps the input signal into the latent space, the approximator $\mathcal{A}$ models a process, and the decoder maps back to the original space. Examples of approximators are conditional generative modeling techniques, e.g., generating an image given a text prompt (condition) [77]. This framework was recently used for 3D shape generation, which are generated in latent space, the final shape in the spatial domain is then constructed by querying the latent representations over a fixed spatial grid [100, 99].
@@ 3 LaM - SLidE
Figure 2: Architecture of our encoder - decoder structure (First Stage): Left: The encoder maps $N$ input tokens to a latent system representation by cross-attending to $L$ learned latent query tokens. The decoder reconstructs the input data from the latent representation using the assigned IDs. Right: Structure of the input token, consisting of an ID, spatial information and features (see also Figure 3).
We introduce an identifier pool and an identifier assignment function which allow us to effectively map and retrieve entities to and from a latent system representation. The ID components preserve the relationships between entities, making them traceable across time-steps. LaM-SLidE follows an encoder $\mathcal{E}$ approximator $\mathcal{A}$ decoder $\mathcal{D}$ paradigm.
@@ 3.1 Problem Formulation
We consider spatial dynamics. Our states $\mathbf{s}\in\mathcal{S}$ describe the configuration of entities within the scene together with their individual features. We assume that a scene consists of $N$ entities $e_{i}$ with $i\in 1,\ldots,N$. An entity $e_{i}$ is described by its spatial location $\mathbf{x}_{i}\in\mathbb{R}^{D_{x}}$ and some further properties $\mathbf{m}_{i}\in\mathbb{R}^{D_{m}}$ (e.g., atom type, etc.), we denote the set of entities as $E=\{e_{1},\dots,e_{n}\}$. We consider states $\mathbf{s}^{t}$ at discretized timepoints $t$. Analogously, we use $\mathbf{x}_{i}^{t}$, $\mathbf{m}_{i}^{t}$ to describe coordinates and properties at time $t$, respectively. We refer to the coordinate concatenation $[\mathbf{x}_{1}^{t},..,\mathbf{x}_{N}^{t}]$ of the $N$ entities in $\mathbf{s}^{t}$ as $\mathbf{X}^{t}\in\mathbb{R}^{N\times D_{x}}$. Analogously, we use $\mathbf{M}^{t}\in\mathbb{R}^{N\times D_{m}}$ to denote $[\mathbf{m}_{1}^{t},..,\mathbf{m}_{N}^{t}]$. When properties are conserved over time, i.e., $\mathbf{M}^{t}=\mathbf{M}^{1}$, we just skip the time index and the time-wise repetition of states and use $\mathbf{M}\in\mathbb{R}^{N\times D_{m}}$. We concatenate sequences of coordinate states $\mathbf{X}^{t}$ with $t\in 1\,..\,T$ to a tensor $\mathbf{X}\in\mathbb{R}^{T\times N\times D_{x}}$, which describes a whole sampled coordinate trajectory of a system with $T$ time points and $N$ entities. An example for such trajectories from dynamical systems are molecular dynamics trajectories (e.g. Figure 11). Notation is summarized in Table 6.
@@ 3.2 Entity Structure Preservation
We aim to preserve the integrity of individual entities in a latent system representation. More specifically, we aim to preserve both the number of entities as well as their structure. For example, in the case of molecules, we want to preserve both the number of atoms and the atom composition. We therefore assign identifiers from an identifier pool to each entity, which allows us to trace the entities by the assigned identifiers. The two key components are: (i) creating a fixed, finite pool of identifiers (IDs) and (ii) defining an unique mapping between entities and identifiers.
@@
For a fixed $u\in\mathbb{N}$, let =${\mathcal{I}}=\{0,1,\dots,u-1\}$ be the identifier pool. An identifier $i$ is an element of the set ${\mathcal{I}}$.
@@
Given an identifier pool ${\mathcal{I}}$ and a finite set of entities $E$, an identifier assignment pool $I$ as defined by Definition 3.2 is non-empty if and only if $|E|\leqslant|{\mathcal{I}}|$.
Proof, see Section F.1.
@@
Given an identifier pool ${\mathcal{I}}$ and a finite set of entities $E$ such that $|E|\leqslant|{\mathcal{I}}|$, the identifier assignment pool $I$ as defined by Definition 3.2 contains finitely many injective functions.
Proof, see Section F.2.
@@
Notably, since $\mathtt{ida(\cdot)}$ may be selected randomly - the specific choice of the mapping $\mathtt{ida}(\cdot)$ can be arbitrary, the only requirement is that an injective mapping between entities and identifiers is established, i.e. each entity is uniquely assigned to an identifier, but not all identifiers need to be assigned to an entity. Further, Proposition 3.3 suggests to use an identifier pool which is large enough, such that a model learned on this identifier pool can generalize across systems with varying numbers of entities.
@@
Aspirin $\text{C}_{9}\text{H}_{8}\text{O}_{4}$ consists of 21 atoms, thus the identifier pool ${\mathcal{I}}$ needs to have at least 21 unique identifiers. We select an assignment function $\mathtt{ida}(\cdot)$, arbitrary, and use it to assign each atom an unique identifier. Notably, e.g. for molecules, we do not explicitly model molecular bond information, as the spatial relationship between atoms (interatomic distances) implicitly capture this information. Figure 3 shows an arbitrary but fixed identifier assignment for aspirin, we illustrate different IDs by colored fingerprint symbols.
@@ 3.3 Model Architecture
Since predicting continuations of system trajectories is a conceptually similar task to generating videos from an initial sequence of images, we took inspiration from Blattmann et al. [15] in using a latent diffusion architecture. We also took inspiration from Jaegle et al. [43] to decompose our model architecture as follows: To map the state of the system composed of $N$ entities to a latent space containing $L$ learned latent tokens ($\in\mathbb{R}^{D_{z}}$), we use a cross-attention mechanism. In the resulting latent space, we aim to train an approximator to predict future latent states based on the embedded initial states. Inversely to the encoder, we again use a cross-attention mechanism to retrieve the physical information of the individual entities of the system from the latent system representation. To wrap it up, LaM-SLidE , is built up by an encoder $\mathcal{E}$ approximator $\mathcal{A}$ decoder $\mathcal{D}$ architecture, $\mathcal{D}\circ\mathcal{A}\circ\mathcal{E}$. A detailed composition of $\mathcal{E}$ and $\mathcal{D}$ is shown in  Figure 2.
@@
The encoder $\mathcal{E}$ aims to encode a state of the system such that the properties of each individual entity $e_{n}$ can be decoded (retrieved) later. At the same time the structure of the latent state representation $\mathbf{Z}^{t}\in\mathbb{R}^{L\times D_{z}}$ is constant and should not depend on the different number $N$ of entities. This contrasts with GNNs, where the number of the latent vectors depends on the number of nodes.
@@
The inputs to the encoder comprise the time dependent location $\mathbf{x}_{n}^{t}\in\mathbb{R}^{D_{x}}$, properties $\mathbf{m}_{n}\in\mathbb{R}^{D_{x}}$, and identity representation $\mathbf{u}_{n}\in\mathbb{R}^{D_{x}}$ of each entity $e_{n}$, as visualized in the right part of Figure 2. We concatenate the different types of features across all entities of the system: $\mathbf{X}^{t}=[\mathbf{x}_{1}^{t},..,\mathbf{x}_{N}^{t}],\;\mathbf{M}=[\mathbf{m}_{1},..,\mathbf{m}_{N}]$ and $\mathbf{U}=[\mathbf{u}_{1},..,\mathbf{u}_{N}]$.
@@
realized by cross-attention [90, 75] between the input tensor $\in\mathbb{R}^{N\times(D_{x}+D_{m}+D_{u})}$, which serves as keys and values, and a fixed number of $L$ learned latent vectors $\in\mathbb{R}^{D_{z}}$ [43], which serve as the queries. The encoding process is depicted on the left side of Figure 2.
@@
As shown in the middle part of Figure 2, $\mathcal{D}$ is realized by cross-attention layers. The latent space representation $\mathbf{Z}^{t}$ serves as the keys and values in the cross-attention mechanism, while the embedded identifier $\mathbf{u}_{n}$ acts as the query. Applied to to all ($n=1,\ldots,N$) system entities $e_{n}$, this results in the retrieved system state information $\mathbf{X}^{t}$ and $\mathbf{M}$. Using the learned identifier embeddings as queries can be interpreted as a form of content-based retrieval and associative memory [6, 40, 75].
@@
Since the generated trajectories should be conditioned on the latent system representations of initial time frames $\mathbf{Z}^{[1\colon T_{o}]}$, we extend $\bm{v}_{\theta}$ with a conditioning argument $\mathbf{C}\in\mathbb{R}^{T\times L\times D_{z}}$, making it effectively a conditional vector field $\bm{v}_{\theta}\colon\mathbb{R}^{T\times L\times D_{z}}\times[0,1]\times\mathbb{R}^{T\times L\times D_{z}}\mapsto\mathbb{R}^{T\times L\times D_{z}}$. The tensor structure of $\mathbf{C}$ is the same as the one for $\mathbf{Z}$. For the first time steps, both tensors have equal values, i.e., $\mathbf{C}^{[1\colon T_{o}]}=\mathbf{Z}^{[1\colon T_{o}]}$. The remaining tensor entries $\mathbf{C}^{[T_{o}+1\colon T]}$ are filled up with mask tokens, see Figure 4. The latent model is structured as series transformer blocks [89, 67], alternating between the spatial and the temporal dimension, see Section B.4. We parametrized our model via a data prediction objective [55], see Section C.1. For architectural details see App. B.

### v4

@@ 1 Introduction
A conventional approach to predict spatial trajectories of entities is to represent scenes as neighborhood graphs and to subsequently process these graphs with graph neural networks (GNNs). When using GNNs [83, 64, 32, 11, 86], each entity is usually represented by a node, and the spatial entities nearby are connected by an edge in the neighborhood graph. Neighborhood graphs have extensively been used for trajectory prediction tasks [51], especially for problems with a large number of indistinguishable entities, [e.g., 81, 63]. Recently, GNNs have been integrated into generative modeling frameworks to effectively capture the behavior of stochastic systems [98, 22].
Despite their widespread use in modeling spatial trajectories, GNNs hardly follow recent trends in latent space modeling, where unified representations [45] together with universality and scalability of transformer blocks [91] offer simple application across datasets and tasks, a behavior commonly observed in computer vision and language processing [24, 25]. Notably, recent breakthroughs in image and video generation can be attributed to latent space generative modeling [38, 15]. In such paradigms, pre-trained encoders and decoders are employed to map data into a latent space, where subsequent modeling is performed, leveraging the efficiency and expressiveness of this representation. This poses the question:
Can we leverage recent techniques from generative latent space modeling to boost the modeling of stochastic trajectories of dynamical systems with a varying number of entities?
@@
Figure 1: Overview of our approach. Left: Conventional graph neural networks (GNNs) model time-evolving systems (e.g., molecular dynamics) by representing entities as nodes and iteratively updating node embeddings and positions to capture system dynamics across timesteps. Right: Latent diffusion models employ an encoder-decoder architecture to compress input data into a lower-dimensional latent space where generative modeling is performed. Latent diffusion models, frequently enhanced with conditional information such as text, excel at generative tasks; however, due to their fixed input/output structure, they are not directly adaptable to physical systems with a varying number of entities. Middle: Our proposed approach LaM-SLidE bridges these paradigms by: (1) introducing identifiers that allow traceability of individual entities, and (2) leveraging a latent system representation.
To leverage a latent system’s representation, we need to be able to trace individual entities within the system. The core idea of LaM-SLidE is the introduction of identifiers (IDs) that allow for the retrieval of entity properties, e.g., entity coordinates, from the latent system representation. Consequently, we can train generative models, like flow-matching [54, 57, 2], purely in latent space, where pre-trained decoder blocks map the generated representations back to the physics domain. An overview is given in Figure˜1. Qualitatively, LaM-SLidE demonstrates flexibility and performs favorably across a variety of different modeling tasks.
@@
  • Cross-domain generalization: We perform experiments in different domains with varying degrees of difficulty, focusing on molecular dynamics, human motion behavior, and particle systems. LaM-SLidE performs favorably with respect to all other architectures and showcases scalability with model size.
@@ 2 Background & Related Work
We note that $\Phi$ does not necessarily need to be defined on the whole space $\mathbb{R}\times\mathcal{S}$, but we assume this for notational simplicity. The exact formal definition of random dynamical systems is more involved and consists of a base flow (noise) and a cocycle dynamical system defined on a physical phase space [8]. We skip the details, but assume that we deal with random dynamical systems for the remainder of the paper. The non-deterministic behavior of such dynamical systems suggests the use of generative modeling approaches.
@@
Recent developments in generative modeling have captured widespread interest. The breakthroughs of the last years were mainly driven by diffusion models [87, 88, 37], a paradigm that transforms a simple distribution into a target data distribution via iterative refinement steps. Flow Matching [54, 57, 2] has emerged as a powerful alternative to diffusion models, enabling simulation-free training between arbitrary start and target distributions [55]. It has also been extended to data manifolds [19]. This approach comes with straighter paths, offering faster integration, and has been successfully applied across different domains like images [27], audio [92], videos [71], protein design [41], and robotics [13].
@@
Latent space modeling has achieved remarkable success at image and video generation [15, 27], where pre-trained encoders and decoders map data into a latent space, and back into the physics space. The latent space aims to preserve the essential structure and features of the original data, often following a compositional structure $\mathcal{D}\circ\mathcal{A}\circ\mathcal{E}$ [85, 4, 5], where the encoder $\mathcal{E}$ maps the input signal into the latent space, the approximator $\mathcal{A}$ models a process, and the decoder maps back to the original space. Examples of approximators include conditional generative modeling techniques, such as generating an image given a text prompt (condition) [77]. This framework was recently used for 3D shape generation, where the final shape in the spatial domain is then constructed by querying the latent representations over a fixed spatial grid [101, 100].
@@ 3 LaM - SLidE
Figure 2: Architecture of our encoder-decoder structure (First Stage): Left: The encoder maps NN input tokens to a latent system representation by cross-attending to $L$ learned latent query tokens. The decoder reconstructs the input data from the latent representation using the assigned IDs. Right: Structure of the input token, consisting of an ID, spatial information, and features (see also Figure˜3).
We introduce an identifier pool and an identifier assignment function which allows us to effectively map and retrieve entities to and from a latent system representation. The ID components preserve the relationships between entities, making them traceable across time-steps. LaM-SLidE follows an encoder $\mathcal{E}$ - approximator $\mathcal{A}$ - decoder $\mathcal{D}$ paradigm.
@@ 3.1 Problem Formulation
We consider spatial dynamics. Our states $\mathbf{s}\in\mathcal{S}$ describe the configuration of entities within the scene together with their individual features. We assume that a scene consists of $N$ entities $e_{i}$ with $i\in 1,\ldots,N$. An entity $e_{i}$ is described by its spatial location $\mathbf{x}_{i}\in\mathbb{R}^{D_{x}}$ and some further properties $\mathbf{m}_{i}\in\mathbb{R}^{D_{m}}$ (e.g., atom type, etc.). We denote the set of entities as $E=\{e_{1},\dots,e_{n}\}$. We consider states $\mathbf{s}^{t}$ at discretized timepoints $t$. Analogously, we use $\mathbf{x}_{i}^{t}$, $\mathbf{m}_{i}^{t}$ to describe coordinates and properties at time $t$, respectively. We refer to the coordinate concatenation $\mathbf{x}_{1}^{t},..,\mathbf{x}_{N}^{t}]$ of the $N$ entities in $\mathbf{s}^{t}$ as $\mathbf{X}^{t}\in\mathbb{R}^{N\times D_{x}}$. Analogously, we use $\mathbf{M}^{t}\in\mathbb{R}^{N\times D_{m}}$ to denote $\mathbf{m}_{1}^{t},..,\mathbf{m}_{N}^{t}]$. When properties are conserved over time, i.e., $\mathbf{M}^{t}=\mathbf{M}^{1}$, we just skip the time index and the time-wise repetition of states and use $\mathbf{M}\in\mathbb{R}^{N\times D_{m}}$. We concatenate sequences of coordinate states $\mathbf{X}^{t}$ with $t\in 1\,..\,T$ to a tensor $\mathbf{X}\in\mathbb{R}^{T\times N\times D_{x}}$, which describes a whole sampled coordinate trajectory of a system with $T$ time points and $N$ entities. An example of such trajectories from dynamical systems are molecular dynamics trajectories (e.g. Figure˜15). Notation is summarized in Table˜6.
@@ 3.2 Entity Structure Preservation
We aim to preserve the integrity of individual entities in a latent system representation. More specifically, we aim to preserve both the number of entities as well as their structure. For example, in the case of molecules, we want to preserve both the number of atoms and the atom composition. We therefore assign identifiers from an identifier pool to each entity, allowing us to trace the entities by their assigned identifiers. The two key components are: (i) creating a fixed, finite pool of identifiers (IDs) and (ii) defining a unique mapping between entities and identifiers.
@@
For a fixed $u\in\mathbb{N}$, let ${\mathcal{I}}=\{0,1,\dots,u-1\}$ be the identifier pool. An identifier $i$ is an element of the set $\mathcal{I}$}.
@@
Given an identifier pool $\mathcal{I}$} and a finite set of entities $E$, an identifier assignment pool $I$ as defined by Definition˜3.2 is non-empty if and only if $|E|\leqslant|{\mathcal{I}}|$.
Proof, see Section˜F.1.
@@
Given an identifier pool $\mathcal{I}$} and a finite set of entities $E$ such that $|E|\leqslant|{\mathcal{I}}|$, the identifier assignment pool $I$ as defined by Definition˜3.2 contains finitely many injective functions.
Proof, see Section˜F.2.
@@
Notably, since $\mathtt{ida(\cdot)}$ may be selected randomly – the specific choice of the mapping $\mathtt{ida}(\cdot)$ can be arbitrary – the only requirement is that an injective mapping between entities and identifiers is established, i.e., each entity is uniquely assigned to an identifier, but not all identifiers need to be assigned to an entity. Further, Proposition˜3.3 suggests using an identifier pool that is large enough, such that a model learned on this identifier pool can generalize across systems with varying numbers of entities.
@@
Aspirin $\text{C}_{9}\text{H}_{8}\text{O}_{4}$ consists of 21 atoms, thus the identifier pool $\mathcal{I}$} needs to have at least 21 unique identifiers. We select an assignment function $\mathtt{ida}(\cdot)$, arbitrary, and use it to assign each atom a unique identifier. Notably, e.g., for molecules, we do not explicitly model molecular bond information, as the spatial relationship between atoms (interatomic distances) implicitly captures this information. Figure˜3 shows an arbitrary but fixed identifier assignment for aspirin; we illustrate different IDs by colored fingerprint symbols.
@@ 3.3 Model Architecture
Since predicting continuations of system trajectories is a conceptually similar task to generating videos from an initial sequence of images, we took inspiration from Blattmann et al. [15] in using a latent diffusion architecture. We also took inspiration from Jaegle et al. [43] to decompose our model architecture as follows: To map the state of the system composed of $N$ entities to a latent space containing $L$ learned latent tokens ($\in\mathbb{R}^{D_{z}}$), we use a cross-attention mechanism. In the resulting latent space, we aim to train an approximator to predict future latent states based on the embedded initial states. Inversely to the encoder, we again use a cross-attention mechanism to retrieve the physical information of the individual entities of the system from the latent system representation. To wrap it up, LaM-SLidE , is built up by an encoder $\mathcal{E}$ - approximator $\mathcal{A}$ - decoder $\mathcal{D}$ architecture, $\mathcal{D}\circ\mathcal{A}\circ\mathcal{E}$. A detailed composition of $\mathcal{E}$ and $\mathcal{D}$ is shown in  Figure˜2.
@@
The encoder $\mathcal{E}$ aims to encode a state of the system such that the properties of each entity $e_{n}$ can be decoded (retrieved) later. At the same time, the structure of the latent state representation $\mathbf{Z}^{t}\in\mathbb{R}^{L\times D_{z}}$ is constant and should not depend on the different number $N$ of entities. This contrasts with GNNs, where the number of latent vectors depends on the number of nodes.
@@
The inputs to the encoder comprise the time dependent location $\mathbf{x}_{n}^{t}\in\mathbb{R}^{D_{x}}$, properties $\mathbf{m}_{n}\in\mathbb{R}^{D_{x}}$, and identity representation $\mathbf{u}_{n}\in\mathbb{R}^{D_{x}}$ of each entity $e_{n}$, as visualized in the right part of Figure˜2. We concatenate the different types of features across all entities of the system: $\mathbf{X}^{t}=[\mathbf{x}_{1}^{t},..,\mathbf{x}_{N}^{t}],\;\mathbf{M}=[\mathbf{m}_{1},..,\mathbf{m}_{N}]$ and $\mathbf{U}=[\mathbf{u}_{1},..,\mathbf{u}_{N}]$.
@@
realized by cross-attention [91, 75] between the input tensor $\in\mathbb{R}^{N\times(D_{x}+D_{m}+D_{u})}$, which serves as keys and values, and a fixed number of $L$ learned latent vectors $\in\mathbb{R}^{D_{z}}$ [43], which serve as the queries. The encoding process is depicted on the left side of Figure˜2.
@@
As shown in the middle part of Figure˜2, $\mathcal{D}$ is realized by cross-attention layers. The latent space representation $\mathbf{Z}^{t}$ serves as the keys and values in the cross-attention mechanism, while the embedded identifier $\mathbf{u}_{n}$ acts as the query. Applied to to all ($n=1,\ldots,N$) system entities $e_{n}$, this results in the retrieved system state information $\mathbf{X}^{t}$ and $\mathbf{M}$. Using the learned identifier embeddings as queries can be interpreted as a form of content-based retrieval and associative memory [6, 40, 75].
@@
Since the generated trajectories should be conditioned on the latent system representations of initial time frames $\mathbf{Z}^{[1\colon T_{o}]}$, we extend $\bm{v}_{\theta}$ with a conditioning argument $\mathbf{C}\in\mathbb{R}^{T\times L\times D_{z}}$, making it effectively a conditional vector field $\bm{v}_{\theta}\colon\mathbb{R}^{T\times L\times D_{z}}\times[0,1]\times\mathbb{R}^{T\times L\times D_{z}}\mapsto\mathbb{R}^{T\times L\times D_{z}}$. The tensor structure of $\mathbf{C}$ is the same as the one for $\mathbf{Z}$. For the first time steps, both tensors have equal values, i.e., $\mathbf{C}^{[1\colon T_{o}]}=\mathbf{Z}^{[1\colon T_{o}]}$. The remaining tensor entries $\mathbf{C}^{[T_{o}+1\colon T]}$ are filled up with mask tokens, see Figure˜4. The latent model is structured as series transformer blocks [90, 67], alternating between the spatial and temporal dimensions (see Section˜B.4). We parametrized our model using a data prediction objective [55], see Section˜C.1. For architectural details see App.˜B.